# SOWA: Adapting Hierarchical Frozen Window Self-Attention to Visual-Language Models for Better Anomaly Detection

### v1

@@ Abstract
Visual anomaly detection is critical in industrial manufacturing, but traditional methods often rely on extensive normal datasets and custom models, limiting scalability. Recent advancements in large-scale visual-language models have significantly improved zero/few-shot anomaly detection. However, these approaches may not fully utilize hierarchical features, potentially missing nuanced details. We introduce a window self-attention mechanism based on the CLIP model, combined with learnable prompts to process multi-level features within a Soldier-Offier Window self-Attention (SOWA) framework. Our method has been tested on five benchmark datasets, demonstrating superior performance by leading in 18 out of 20 metrics compared to existing state-of-the-art techniques.
@@ 1 Introduction
Visual anomaly classification (AC) and segmentation (AS) are crucial tasks in industrial manufacturing, aimed at identifying and localizing defects within images [2, 4, 30, 24, 3, 36]. The main challenge is the rarity and variability of defects, leading to a scarcity of representative anomaly samples in training data. Most existing methods focus on one-class or unsupervised anomaly detection, requiring numerous normal images and custom models for each task, making them not scalable across many tasks [30, 11, 4, 9, 10, 19, 29, 39, 44]. While achieving high accuracy with large normal datasets [30, 5, 10], their performance drops significantly with fewer normal samples [15, 32, 44], highlighting the need for improved methods.
Vision-language models (VLMs) [27, 17, 13], such as CLIP, have shown promise in zero/few-shot classification tasks by using large-scale training on vision-language annotated pairs via contrastive learning [7], enabling them to capture broad concepts and generalize to unseen tasks. For example, WinCLIP [16] is built on the CLIP model and uses a multi-scale window-based strategy for segmentation. By introducing text prompts to guide anomaly detection, WinCLIP effectively performs zero/few-shot anomaly detection without needing extensive training on each specific category. Similarly, April-GAN [8] introduces extra linear layers to map image features from the CLIP image encoder to the text feature space, enabling precise similarity comparisons and resulting in detailed anomaly maps. The main advantage of CLIP-based approaches is their strong generalization capabilities due to large-scale vision-language pretraining. By using language to provide context, these models can better distinguish between normal and anomalous states, even with limited normal samples, which is particularly useful when obtaining anomaly samples is difficult.
However, these approaches also have limitations. Firstly, they often do not fully leverage the hierarchical features captured by the CLIP vision transformer (ViT) [12], potentially missing nuanced details necessary for accurate anomaly detection. Secondly, although effective at leveraging text prompts, the fixed encoded prompts are originally designed for final global features and may not optimally balance local features from shallow layers. These limitations can lead to challenges in detecting anomalies that vary widely in scale and context.
@@
  • A elaborately designed framework is proposed, leveraging the power of vision-language models for anomaly detection.
  • The method is tested on several benchmark datasets, demonstrating superior performance compared to existing state-of-the-art techniques in both anomaly classification and segmentation tasks.
  • The working mechanism of the model is validated, offering valuable insights for future research.
@@ 2 Related Work
Figure 2: (a) The architecture of Hierarchical Frozen Window Self-Attention (HFWA) model; (b) Detailed structure of FWA (Frozen Window Attention) adapter; (c) Detailed structure of FWA.
@@
This work focuses on anomaly detection (AD). We define a training dataset $D_{train}$ with annotated samples from the target domain, where each sample is a tuple $\{(x_{i},c_{i},S_{i})\}$: $x_{i}$ is the training image, $c_{i}\in\{-1,+1\}$ is the image-level anomaly classification label, and $S_{i}\in\{-1,+1\}^{h\times w}$ is the pixel-level anomaly segmentation mask for images of size $h\times w$. ’+1’ indicates anomaly, and ’-1’ indicates normal. The model $M_{adapt}$ aims to predict both image-level ($c$) and pixel-level anomalies ($S$). Image-level prediction is a binary classification indicating if anomalies exist, while pixel-level prediction identifies anomaly locations within the image. For zero-shot learning, the test dataset $D_{test}$ contains data from different domains than $D_{train}$ to evaluate generalization. For few-shot learning, a reference set $T^{\text{r}ef}=\{D_{\alpha}^{\text{r}ef}\}$ contains small reference datasets from the same domain as test samples. For each test sample $x_{\text{t}est}$ in $D_{test}$, $M_{\text{a}dapt}$ uses $K$ randomly selected reference images from the corresponding $D_{\alpha}^{\text{r}ef}$ in $T^{\text{r}ef}$ to evaluate $x_{\text{t}est}$.
@@ 5.2 Main Results
Metric Dataset WinCLIP VAND Ours
@@
Figure 5: Comparisons with few-shot (K=4) anomaly detection methods on datasets of MVTec-AD, Visa, BTAD, DAGM and DTD Synthetic.
@@
Source (Venue) MVTec-AD Visa
@@
Quantitative Results. The results, as shown in Table 1, demonstrate that our proposed method consistently outperforms WinCLIP and VAND across multiple datasets and evaluation metrics. Our method consistently demonstrated superior performance across the datasets. On the MVTec-AD dataset, our method achieved the highest AC AUROC score of 96.8±0.3, surpassing WinCLIP by 1.6 percentage points and VAND by 4.0 percentage points. Similarly, in the Visa dataset, our method achieved an AC AUROC score of 92.9±0.2, exceeding WinCLIP by 5.6 percentage points and VAND by 0.3 percentage points. On the BTAD dataset, our method reached 94.8±0.2, which is 7.8 percentage points higher than WinCLIP and 2.7 percentage points higher than VAND. In the DAGM dataset, our method achieved an impressive score of 99.1, outperforming WinCLIP by 5.3 percentage points and VAND by 4.5 percentage points. Lastly, on the DTD-Synthetic dataset, our method scored 99.1±0.0, which is 1.0 percentage point higher than WinCLIP and 0.6 percentage points higher than VAND. These results highlight the robustness and accuracy of our approach in image-level anomaly detection.
In terms of pixel-level performance, our method also showed significant improvements. Across all datasets, our method demonstrated superior performance in AS PRO, consistently leading over both WinCLIP and VAND. For instance, on the DAGM dataset, our method achieved an exceptional AS PRO score of 96.6, surpassing WinCLIP by 14.2 percentage points and VAND by 20.1 percentage points. This demonstrates our method’s effectiveness in detecting fine-grained anomalies at the pixel level.
@@ Appendix E Zero-shot Performance Comparison
Metric Dataset WinCLIP VAND AnormalyCLIP Ours
@@
In the zero-shot setting, our analysis focuses on the performance of various models across different datasets for four key metrics: AC AUROC, AC AP, AS AUROC, and AS AUPRO. The models compared are WinCLIP, VAND, AnormalyCLIP [42], and our proposed model. The results are shown in Table 5. For AC AUROC, AnormalyCLIP consistently outperforms the other models on most datasets, notably achieving the highest scores on BTAD (88.3), DAGM (97.5), and Visa (82.1). Our model demonstrates competitive performance on DAGM (96.7) and DTD-Synthetic (95.2), but lags behind AnormalyCLIP on several datasets.

### v3

@@ Abstract
Visual anomaly detection is essential in industrial manufacturing, yet traditional methods often rely heavily on extensive normal datasets and task-specific models, limiting their scalability. Recent advancements in large-scale vision-language models have significantly enhanced zero- and few-shot anomaly detection. However, these approaches may not fully leverage hierarchical features, potentially overlooking nuanced details crucial for accurate detection. To address this, we introduce a novel window self-attention mechanism based on the CLIP model, augmented with learnable prompts to process multi-level features within a Soldier-Officer Window Self-Attention (SOWA) framework. Our method has been rigorously evaluated on five benchmark datasets, achieving superior performance by leading in 18 out of 20 metrics, setting a new standard against existing state-of-the-art techniques.
@@ 1 Introduction
Visual anomaly classification (AC) and segmentation (AS) are crucial tasks in industrial manufacturing, aimed at identifying and localizing defects within images [2, 4, 30, 24, 3, 36]. The primary challenge of these tasks is the rarity and variability of defects, leading to a scarcity of representative anomaly samples in training data. Most existing methods emphasize one-class or unsupervised anomaly detection, requiring a large volume of normal images and task-specific models, which limits scalability across diverse tasks [30, 11, 4, 9, 10, 19, 29, 39, 44]. Although high accuracy can be achieved with ample normal data [30, 5, 10], performance declines significantly with fewer normal samples [15, 32, 44], making data scarcity a key bottleneck for improved performance. To reduce reliance on extensive data, we need to seek more robust and advanced methods to tackle this problem.
Vision-language models (VLMs) [27, 17, 13], such as CLIP, have demonstrated strong performance in zero- and few-shot classification tasks through large-scale training on vision-language annotated pairs via contrastive learning [7]. This training enables VLMs to capture broad concepts and generalize to new tasks. For instance, WinCLIP [16], an extension of CLIP, applies a multi-scale window-based strategy for segmentation, using text prompts to guide anomaly detection and effectively performing zero- and few-shot anomaly detection without extensive category-specific training. Similarly, April-GAN [8] introduces additional linear layers to map image features obtained from CLIP’s image encoder into the text feature space, facilitating precise similarity comparisons and generating detailed anomaly maps. The main advantage of CLIP-based approaches lie in their strong generalization capabilities, achieved through large-scale vision-language pretraining. By leveraging language to provide contextual understanding, these models can effectively differentiate between normal and anomalous states, even with limited normal samples, which is particularly valuable when anomaly samples are difficult to obtain.
However, these approaches also have limitations. First, they often underutilize the hierarchical features captured by the CLIP vision transformer (ViT) [12], potentially missing nuanced details crucial for accurate anomaly detection. Second, although effective in using text prompts, the fixed encoded prompts are primarily designed for final global features, which may not optimally integrate local features from earlier layers. These limitations can create challenges in detecting anomalies that vary significantly in scale and context.
@@
  • We propose a carefully crafted framework that leverages the strengths of vision-language models for anomaly detection.
  • The method is evaluated on multiple benchmark datasets, demonstrating superior performance compared to existing state-of-the-art techniques in both anomaly classification and segmentation tasks.
  • The model’s operational mechanism is validated, providing valuable insights that can inform future research.
@@ 2 Related Work
Figure 2: (a) The architecture of Hierarchical Frozen Window Self-Attention (HFWA) model. Block H1 to H4 represent the four feature extraction stages of CLIP ViT; (b) Detailed structure of FWA (Frozen Window Attention) adapter; (c) Detailed structure of FWA.
@@
Let $D_{train}=\{(x_{i},c_{i},S_{i})\}$ denote a dataset containing annotated samples. Each sample is given as a tuple $(x_{i},c_{i},S_{i})$, where $x_{i}$ is an image of size $h\times w$, $c_{i}\in\{-1,+1\}$ represents the image-level anomaly classification label, and $S_{i}\in\{-1,+1\}^{h\times w}$ is a pixel-level anomaly segmentation mask of size $h\times w$, with ‘1’ indicating anomaly, and ‘-1’ indicating normal. Given an image $x$, we aim to predict both image-level ($c$) and pixel-level anomalies ($S$). The image-level prediction is a binary classification that indicates whether anomalies are present, while the pixel-level prediction pinpoints the specific locations of anomalies within the image. In zero-shot learning, the test dataset includes data from domains distinct from the training dataset to assess the model’s generalization ability. For few-shot learning, a reference set $T^{\text{r}ef}=\{D_{\alpha}^{\text{r}ef}\}$ contains small reference datasets from the same domain as test samples. For each test sample $x_{\text{t}est}$ in $D_{test}$, $M_{\text{a}dapt}$ uses $K$ randomly selected reference images from the corresponding $D_{\alpha}^{\text{r}ef}$ in $T^{\text{r}ef}$ to evaluate $x_{\text{t}est}$.
@@ 5.2 Main Results
Metric Dataset WinCLIP April-GAN Ours
@@
Figure 5: Comparisons with zero-/few-shot anomaly detection methods on datasets of MVTec-AD, Visa, BTAD, DAGM and DTD Synthetic.
@@
Method (Source) MVTec-AD Visa
@@
Quantitative Results. The results, as shown in Table 1, demonstrate that our proposed method consistently outperforms WinCLIP and April-GAN across multiple datasets and evaluation metrics. Our method consistently demonstrated superior performance across the datasets. On the MVTec-AD dataset, our method achieved the highest AC AUROC score of 96.8±0.3, surpassing WinCLIP by 1.6 percentage points and April-GAN by 4.0 percentage points. Similarly, in the Visa dataset, our method achieved an AC AUROC score of 92.9±0.2, exceeding WinCLIP by 5.6 percentage points and April-GAN by 0.3 percentage points. On the BTAD dataset, our method reached 94.8±0.2, which is 7.8 percentage points higher than WinCLIP and 2.7 percentage points higher than April-GAN. In the DAGM dataset, our method achieved an impressive score of 99.1, outperforming WinCLIP by 5.3 percentage points and April-GAN by 4.5 percentage points. Lastly, on the DTD-Synthetic dataset, our method scored 99.1±0.0, which is 1.0 percentage point higher than WinCLIP and 0.6 percentage points higher than April-GAN. These results highlight the robustness and accuracy of our approach in image-level anomaly detection.
In terms of pixel-level performance, our method also showed significant improvements. Across all datasets, our method demonstrated superior performance in AS PRO, consistently leading over both WinCLIP and April-GAN. For instance, on the DAGM dataset, our method achieved an exceptional AS PRO score of 96.6, surpassing WinCLIP by 14.2 percentage points and April-GAN by 20.1 percentage points. This demonstrates our method’s effectiveness in detecting fine-grained anomalies at the pixel level.
@@ Appendix E Zero-shot Performance Comparison
Metric Dataset WinCLIP April-GAN AnormalyCLIP Ours
@@
In the zero-shot setting, our analysis focuses on the performance of various models across different datasets for four key metrics: AC AUROC, AC AP, AS AUROC, and AS AUPRO. The models compared are WinCLIP, April-GAN, AnormalyCLIP [42], and our proposed model. The results are shown in Table 5. For AC AUROC, AnormalyCLIP consistently outperforms the other models on most datasets, notably achieving the highest scores on BTAD (88.3), DAGM (97.5), and Visa (82.1). Our model demonstrates competitive performance on DAGM (96.7) and DTD-Synthetic (95.2), but lags behind AnormalyCLIP on several datasets.